OpenStudy (anonymous):

Part 1: Solve the inequality −5x + 8 less than or greater to x − 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

On Part 1, did you mean "less than or equal to x - 4"?

OpenStudy (anonymous):

@jazy Nope . Thats What My Question Said

OpenStudy (anonymous):

idk how to do this wat grade is it

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

Algebra?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i hate algebra :(

OpenStudy (anonymous):

Me Too !

OpenStudy (anonymous):

It wasn't that bad for me after I got the hang of it. Are you SURE it doesn't say, Solve the inequality −5x + 8 (Greater Then or equal to) x − 4 and show all of your work. You can use <= to represent the less than or equal to symbol.

OpenStudy (anonymous):

No But I Will Try It Because Maybe My Teacher Mad A Mistake

OpenStudy (anonymous):

I mean't *Less Than or equal to

OpenStudy (anonymous):

So Can You Help Me ?

OpenStudy (anonymous):

Yep(: Solve the inequality −5x + 8 less than or equal to x − 4 and show all of your work. You can use <= to represent the less than or equal to symbol. -5x + 8 <= x - 4 Combine like terms. -5x + 8 <= x - 4 Subtract 8 from both sides. -5x <= x - 12 Add x to both sides. -4x <= -12 Divide both sides by -4. Flip the sign to >= since we are divide by a negative(-) number. x >= (-12 / -4) What is -12 divided by -4?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yep! So: x >= 3

OpenStudy (anonymous):

To graph keep in mind that: Greater/Less than or EQUAL to(>= or <=)... Closed circle Greater/Less than(> or <)... Open circle Greater than(> or >=)... Shading to the RIGHT Less than( < or <=)... Shading to the LEFT

OpenStudy (anonymous):

Thank You Soo Much (:

OpenStudy (anonymous):

No problem! Glad I could help. (: